OpenStudy (anonymous):

The displacement d (in meters) of an object at time t (in seconds) is given. d= -sin(t/9) (a) Describe the motion of the object. is it simple harmonic motion or damped motion? (b) What is the maximum displacement from its resting position? Round your answer to two decimal places. in meters (c) What is the time required for one oscillation? Round your answer to two decimal places. in seconds (d) What is the frequency? Round your answer to two decimal places. in oscillation/second

OpenStudy (amistre64):

simple harmonic i believe; there is no part of it that is accounting for a dampned effect

OpenStudy (anonymous):

1 is the max displacement

OpenStudy (amistre64):

the displacement from rest is the amplitude; which is 1 the initial displacement tho is -1

OpenStudy (anonymous):

it asks for decimal places tho?

OpenStudy (anonymous):

initial disp is 0 amistre

OpenStudy (amistre64):

1.0000 then lol

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

hahaha

OpenStudy (amistre64):

well.... yeah, initial = 0 :)

OpenStudy (anonymous):

1/18pi is the frequency

OpenStudy (anonymous):

and hence18pi is the time period for one oscillation

OpenStudy (anonymous):

so max disp is 0 or 1?

OpenStudy (amistre64):

1

OpenStudy (amistre64):

are you accounting for the pi already in the problem? or are you atributing an external pi?

OpenStudy (amistre64):

sin[(1/9)t]

OpenStudy (anonymous):

see the eqn for shm is x=sin (wt)

OpenStudy (anonymous):

and 2(pi)f=w

OpenStudy (anonymous):

lol i got lost

OpenStudy (anonymous):

so f=w/2pi

OpenStudy (anonymous):

w=1/9 hence f = 1/18pi

OpenStudy (amistre64):

i get turned around whe w = pi/n lol

OpenStudy (anonymous):

and since T=1/f , T=18pi

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=-sin%28t%2F9%29 hims good on it

OpenStudy (anonymous):

yeah waves were my forte ;)

OpenStudy (anonymous):

so was c) 18pi ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did u follow the working?

OpenStudy (anonymous):

lol got lost

OpenStudy (anonymous):

would you mind explaining it to me?

OpenStudy (anonymous):

do u know what simple harmonic motion is?

OpenStudy (anonymous):

i did it before but forgot. is it got to do with the pendulums?

OpenStudy (anonymous):

yeah...anything in shm follows the displacement eqn x = Asin(wt) w is called the angular frequency and equals 2pi x f where f is the frequency

OpenStudy (anonymous):

and f=1/T where T is the time for one cycle or oscillation

OpenStudy (anonymous):

i remember now haha

OpenStudy (anonymous):

max disp from its resting position isnt 1 tho

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

im not sure. but the answer u got is 1.00? its not right for some reason

OpenStudy (anonymous):

so its -1?

OpenStudy (anonymous):

is it -1.00??

OpenStudy (anonymous):

must be that then...

OpenStudy (anonymous):

no it isn't.

OpenStudy (anonymous):

it must be to two dp and be in meters

OpenStudy (anonymous):

gotta be 1.00...dunno y its coming out wrong..

OpenStudy (anonymous):

me either but it aint working. tried it already

OpenStudy (anonymous):

cud it be 0

OpenStudy (anonymous):

how do u work that part anyways?

OpenStudy (anonymous):

hey amistre, the max disp from rest is turnin out wrong

OpenStudy (anonymous):

i put in -1.00 and 1.00 both r wrong

OpenStudy (amistre64):

thats becasue we assumed displacement = amplitude d = displacement; so we gotta work the formula :)

OpenStudy (anonymous):

amplitude is defined as the max displacemnt from the mean position

OpenStudy (amistre64):

it peaks at |+-1| it aint asking for the value of 't' at the max becasue there aint no interval to associate a single t into

OpenStudy (amistre64):

have you tried it without the decimals?

OpenStudy (amistre64):

just enter: 1

OpenStudy (anonymous):

its my last shot tho lol

OpenStudy (anonymous):

nah, it didnt work